Participants: E. Adli, F. Antoniou, H. Burkhardt, B. Dalena, E. Marin, M. Martini, Ö. Mete, G. Morpurgo, Y. Papaphilippou, L. Rinolfi, D. Schulte, F. Stulle, J. Snuverink, T. Touze, A. Vivoli
D. Schulte reported on discussions on budget and fellow positions at the CSC meeting. For the ACE meeting a description has to be prepared of the work which will become part of the CDR. A draft was sent for review by D.S. to all people involved. Next Monday D.S. will organize a meeting of all speakers reporting on beam dynamics at the ACE meeting. For the ILC CLIC LET Beam Dynamics Workshop in June D.S. asked for proposals of talks. The registration is now open.
Short updates on current work:
F. Stulle reviewed some limitations of the current drive beam feedforward. A lot more flexibility in the setup of the bc system is allowed without spoiling the RF creation in the PETS. He also performed simulations of the main beam RTML using 500GeV parameters. A major revision required due to the higher charge.
A. Vivoli reported that he has implemented the algorithms of the IBS code Mocac into a C program. First comparisons are encouraging but more work is required.
F. Antoniou is currently fixing the lattice for the pre-damping rings.
Ö. Mete finished the analysis of the latest measurement data acquired at PHIN. Unfortunately, the accuracy of the beam size measurements are limited by the instrumentation.
H. Burkhardt informed that halo and tail simulations for the drive beam gave auseful result and were presented as a poster at PAC09. They must be continued and simulations must be improved. To improve the models in GEANT4 measurements at beam energies from 100MeV to 1GeV would be interesting.
B Dalena improved simulations on the impact of the detector solenoid field on beam dynamics.
Y. Papaphilippou revised the damping rings. Their layout is close to being frozen. The work for 500GeV parameters is ongoing. A prototype of the superconducting wiggler is ready for being tested. Coated chambers has been put into CESR-TA for testing.
T. Touze started to write a new code for pre-alignment simulations since the existing one is too limited.
E. Adli finished two reports on PETS test and is performing simulations for the ACE meeting.
D. Schulte had been at meetings of review committees for ILC and SuperB.
L. Rinolfi is working on the baseline configuration for the main beam injectors. A discussion is going on to produce CLIC-like bunches at the SLAC gun test stand. LAL started working on the positron source.
M. Martini is working on IBS and reviewed the algorithms of Mocac (see slides).
There are minutes attached to this event.
Show them.